OpenStudy (anonymous):

10-8x=-2x+4x ?? help

OpenStudy (solomonzelman):

Check the problem carefully. I don't think you really meant what you wrote. But if you DID write what you meant, then.... 1) Add 8x to both sides 2) add the Xs on the right side 3) Divide both sides by 10.

OpenStudy (anonymous):

yes you got it thank you

OpenStudy (solomonzelman):

you welcome.

OpenStudy (anonymous):

so it would be -5?

OpenStudy (solomonzelman):

10-8x=-2x+4x +8x +8x 10=-2x+4x+8x 10=(-2+4+8)x 10=(-10)x can you tell me the answer ?

OpenStudy (anonymous):

-1?

OpenStudy (solomonzelman):

Oh yeah, you got the correct answer !!

OpenStudy (anonymous):

ya thanks i suck at these